# Joint Learning of Linear Dynamical Systems under Smoothness Constraints

### v1

@@ Abstract
We consider the problem of joint learning of multiple linear dynamical systems. This has received significant attention recently under different types of assumptions on the model parameters. The setting we consider involves a collection of $m$ linear systems each of which resides on a node of a given undirected graph $G=([m],\mathcal{E})$. We assume that the system matrices are marginally stable, and satisfy a smoothness constraint w.r.t $G$ – akin to the quadratic variation of a signal on a graph. Given access to the states of the nodes over $T$ time points, we then propose two estimators for joint estimation of the system matrices, along with non-asymptotic error bounds on the mean-squared error (MSE). In particular, we show conditions under which the MSE converges to zero as $m$ increases, typically polynomially fast w.r.t $m$. The results hold under mild (i.e., $T\sim\log m$), or sometimes, even no assumption on $T$ (i.e. $T\geq 2$).
@@ 1 Introduction
with $A^{*}$ denoting the unknown system matrix to be estimated, and $(\eta_{t})_{t\geq 0}$ denotes the (unobserved) process noise – typically assumed to be centered and independent over $t$. While the problem has been studied extensively in the literature, classical work on linear system identification typically either consisted of asymptotic guarantees (e.g., (Lai and Wei, 1982, 1983, 1986)), or had sub-optimal bounds that are exponential in the degree of the linear system (Campi and Weyer, 2002; Vidyasagar and Karandikar, 2008).
@@ 1.1 Finite-time identification of LDSs
A recent line of work has focused on finite-time identification of linear systems which amounts to deriving non-asymptotic error bounds for estimating $A^{*}$ (e.g., (Shirani Faradonbeh et al., 2018; Simchowitz et al., 2018; Sarkar and Rakhlin, 2019; Jedra and Proutiere, 2020)). These results provide error bounds for estimating $A^{*}$ via the ordinary least-squares (OLS) estimator (yielding $\widehat{A}$), with the error quantified using the spectral norm $\big{\|}{\widehat{A}-A^{*}}\big{\|}_{2}$. The error is shown to decay to zero at the (near-) parametric rate [^1] Throughout, $\tilde{O}(\cdot)$ hides logarithmic factors. $\tilde{O}(1/\sqrt{T})$ provided $T$ is suitably large enough, and under different assumptions on the spectral radius [^2] See Section 2.1 for a description of notation. of $A^{*}$, namely $\rho(A^{*})$. A typical assumption in this regard is that of stability, which amounts to saying that either $\rho(A^{*})<1$ (strict stability), or even $\rho(A^{*})\leq 1$ (marginal stability). The state-of-art result in this regard is arguably that of Sarkar and Rakhlin (2019) – when $A^{*}$ is marginally stable, it is shown that with probability at least $1-\delta$,
$\big{\|}{\widehat{A}-A^{*}}\big{\|}_{2}\lesssim\sqrt{\frac{d\log(\frac{1+\operatorname{Tr}(\Gamma_{T}(A^{*}))}{\delta})}{T}}\quad\text{ if }\quad T\gtrsim d\log\Big{(}\frac{1+\operatorname{Tr}(\Gamma_{T}(A^{*}))}{\delta}\Big{)}.$ (1.2)
@@ 1.2 Learning multiple LDSs under smoothness constraints
which means that the quadratic variation of $A^{*}_{1},\dots,A^{*}_{m}$ w.r.t $G$ is bounded. For bounded $A^{*}_{l}$’s one can obviously find $S_{m}=O(m)$ such that (1.5) is true, however the interesting smoothness regime is when $S_{m}=o(m)$. Indeed, in this regime, the node-dynamics of the graph $G$ share a non-negligible amount of information between themselves, and thus one might hope to leverage this fact in order to jointly estimate $A^{*}_{l}$’s meaningfully. In this regard, the following “sanity checks” are useful.
@@
We propose two estimators for this problem, and derive non-asymptotic error bounds on the mean-squared error (MSE) $\frac{1}{m}\sum_{l=1}^{m}\big{\|}{\widehat{A}_{l}-A^{*}_{l}}\big{\|}_{F}^{2}$, where $\widehat{A}_{l}$ is the estimate of $A^{*}_{l}$.
@@ 2.1 Notation
For any vector $x\in\mathbb{R}^{n}$, $\big{\|}{x}\big{\|}_{p}$ denotes the usual $\ell_{p}$ norm of $x$. For $X\in\mathbb{R}^{n\times m}$, we denote $\big{\|}{X}\big{\|}_{2}$, $\big{\|}{X}\big{\|}_{F}$ to be respectively the spectral and Frobenius norms of $X$, while $\Big{\langle}X,Y\Big{\rangle}=\operatorname{Tr}(X^{\top}Y)$ denotes the inner product between $X$ and $Y$ (with $\operatorname{Tr}(\cdot)$ denoting the trace operator). For $n\times n$ matrices X with eigenvalues $\lambda_{i}\in\mathbb{C}$, we denote $\rho(X):=\max_{i=1,\dots,n}\left|{\lambda_{i}}\right|$ to be the spectral radius of $X$. Also, $\kappa(X)(\geq 1)$ denotes the condition number of $X$ (ratio of largest and smallest singular values). For a positive definite matrix $B$, and a vector $x$, we will denote
@@
Finally, we denote $\mathbf{1}_{\mathcal{B}}$ to be the indicator function for an event $\mathcal{B}$.
@@ 2.2 Setup
$\sum_{\left\{{l,l^{\prime}}\right\}\in\mathcal{E}}\big{\|}{A^{*}_{l}-A^{*}_{l^{\prime}}}\big{\|}_{F}^{2}=a^{*}(L\otimes I_{d^{2}})a^{*}\leq S_{m}.$ (2.1)
@@ 2.3 Laplacian smoothing
The first is the penalized least-squares (LS) method
$(\widehat{A}_{1},\dots,\widehat{A}_{m})=\underset{A_{1},\dots,A_{m}\in\mathbb{R}^{d\times d}}{\operatorname{argmin}}\ \left(\sum_{i=1}^{m}\big{\|}{x_{l,{t+1}}-A_{l}x_{l,t}}\big{\|}_{F}^{2}+\lambda\sum_{\left\{{l,l^{\prime}}\right\}\in\mathcal{E}}\big{\|}{A_{l}-A_{l^{\prime}}}\big{\|}_{F}^{2}\right)$ (2.3)
@@ 2.4 Subspace-constrained LS
Denote $P^{(\tau)}=\sum_{l=m-\tau+1}^{m}v_{l}v_{l}^{\top}$ to be the projection matrix for the low-frequency part of $L$, we obtain another projection matrix $\widetilde{P}^{(\tau)}=P^{(\tau)}\otimes I_{d^{2}}$. The condition (2.1) implies that $a^{*}\in\mathbb{R}^{md^{2}}$ – formed by column-wise stacking of $\operatorname{vec}(A^{*}_{1}),\dots,\operatorname{vec}(A^{*}_{m})$ – lies close to the low frequency part of $L\otimes I_{d^{2}}$. Hence, $a^{*}\approx\widetilde{P}^{(\tau)}a^{*}$ for a suitable $\tau$ depending on $S_{m}$. For example, if $S_{m}=0$, then $\tau=1$ and clearly $a^{*}=\widetilde{P}^{(\tau)}a^{*}$.
@@ 3.1 Laplacian smoothing
The proof is motivated by ideas in the proof of (Sarkar and Rakhlin, 2019, Theorem 1) for learning marginally stable linear dynamical systems. The first error term in (1) corresponds to the bias, while the remaining two terms therein constitute the variance. The condition on $T$ in (3.1) is an artefact of the analysis, and arises from Lemma 1 in order to ensure $Q^{\top}Q+\lambda(L\otimes I_{d^{2}})\succ 0$. In general, obtaining lower bounds on the eigenvalues of $Q^{\top}Q+\lambda(L\otimes I_{d^{2}})$ is challenging as the eigenspaces of $L\otimes I_{d^{2}}$ and $Q^{\top}Q$ are not aligned. Our strategy is the admittedly naive approach that instead seeks to ensure $Q^{\top}Q\succ 0$ using results from (Sarkar and Rakhlin, 2019, Section 9).
@@
    Hence if $S_{m}=O(m^{\epsilon})$ for some constant $\epsilon\in(0,1)$, then we require $m=\Omega(\operatorname{poly}(d,T^{\alpha}))$ to drive the MSE below a given threshold. Recall that $\alpha\in[2d]$, and depends on the size of the largest Jordan blocks of $A^{*}_{l}$’s. If for each $l$, the largest Jordan block of $A^{*}_{l}$ is of constant size, then $\alpha\asymp 1$. But in case $\alpha\asymp d$, then $m$ will grow exponentially with $d$ in our bounds. This exponential dependence on $\alpha$ is essentially a consequence of the $\gamma_{1}(T,m,\delta)$ term appearing outside the logarithm in the third term of (1), and is again an artefact of the analysis.
@@
Suppose each $A^{*}_{l}$ is marginally stable, i.e. satisfies (3.4), and recall $\alpha,\kappa_{\max}$ in (3.6). Under Assumptions 1 and 2, suppose $\eta_{l,t}$ has $R$-subgaussian entries for each $l,t$.
@@
    $\tilde{S}_{m}:=S_{m}+m^{2}p_{1}(a^{*});\quad\sum_{i=1}^{d^{2}}\Big{\langle}a^{*},v_{1}\otimes e_{i}\Big{\rangle}^{2}\leq p_{1}(a^{*}).$
@@
As a sanity check, let us verify that we obtain better bounds than what we would have by naively estimating each $A^{*}_{l}$ individually. Indeed, assuming $R$ to be a constant, if (3.4) holds, then applying the OLS for each node $l$, we obtain using (1.2) and a simple union bound that with probability at least $1-\delta$,
$\frac{1}{m}\sum_{l=1}^{m}\big{\|}{\widehat{A}-A^{*}_{l}}\big{\|}_{F}^{2}\lesssim\frac{d^{2}}{mT}\sum_{l=1}^{m}\log\Big{(}\frac{m(1+\operatorname{Tr}(\Gamma_{T}(A^{*}_{l})))}{\delta}\Big{)}\lesssim\frac{d^{2}}{T}\max_{l\in[m]}\log\Big{(}\frac{m(1+\operatorname{Tr}(\Gamma_{T}(A^{*}_{l})))}{\delta}\Big{)},$ (3.12)
provided $T\gtrsim d\max_{l\in[m]}\log(\frac{m[1+\operatorname{Tr}(\Gamma_{T}(A^{*}_{l}))]}{\delta})$. This condition on $T$ is essentially the same as in (3.1), and is ensured for $T=\Omega(\log m)$. However in the regime $T=\Theta(\log m)$, the bound in (3.12) is $\Omega(1)$ w.r.t $m$.

### v2

@@ Abstract
We consider the problem of joint learning of multiple linear dynamical systems. This has received significant attention recently under different types of assumptions on the model parameters. The setting we consider involves a collection of $m$ linear systems, each of which resides on a node of a given undirected graph $G=([m],\mathcal{E})$. We assume that the system matrices are marginally stable, and satisfy a smoothness constraint w.r.t $G$ – the smoothness measured in a manner akin to the quadratic variation of a signal on a graph. Given access to the states of the nodes over $T$ time points, we then propose two estimators for joint estimation of the system matrices, along with non-asymptotic error bounds on the mean-squared error (MSE). In particular, we show conditions under which the MSE converges to zero as $m$ increases, typically polynomially fast w.r.t $m$. The results hold under mild (i.e., $T\sim\log m$), or sometimes, even no assumption on $T$ (i.e. $T\geq 2$).
@@ 1 Introduction
with $A^{*}$ denoting the unknown system matrix to be estimated, and $(\eta_{t})_{t\geq 1}$ denotes the (unobserved) process noise – typically assumed to be centered and independent over $t$. While the problem has been studied extensively in the literature, classical work on linear system identification typically either consisted of asymptotic guarantees on the estimation error (e.g., (Lai and Wei, 1982, 1983, 1986)), or consisted of quantities which are exponential in the degree of the linear system (Campi and Weyer, 2002; Vidyasagar and Karandikar, 2008).
@@ 1.1 Finite-time identification of LDSs
A recent line of work has focused on finite-time identification of linear systems, which amounts to deriving non-asymptotic error bounds for estimating $A^{*}$ (e.g., (Shirani Faradonbeh et al., 2018; Simchowitz et al., 2018; Sarkar and Rakhlin, 2019; Jedra and Proutiere, 2020)). These results provide error bounds for estimating $A^{*}$ via the ordinary least-squares (OLS) estimator (yielding $\widehat{A}$), with the error quantified using the spectral norm $\big{\|}{\widehat{A}-A^{*}}\big{\|}_{2}$. The error is shown to decay to zero at the (near-) parametric rate [^2] Throughout, $\tilde{O}(\cdot)$ hides logarithmic factors. $\tilde{O}(1/\sqrt{T})$ provided $T$ is suitably large enough, and under different assumptions on the spectral radius [^3] See Section 2.1 for a description of notation. $\rho(A^{*})$. A typical assumption in this regard is that of stability, which amounts to saying that either $\rho(A^{*})<1$ (strict stability), or even $\rho(A^{*})\leq 1$ (marginal stability). The state-of-art result in this regard is arguably that of Sarkar and Rakhlin (2019) – when $A^{*}$ is marginally stable, it is shown that with probability at least $1-\delta$,
$\big{\|}{\widehat{A}-A^{*}}\big{\|}_{2}\lesssim\sqrt{\frac{d\log\left(\frac{1+\operatorname{Tr}(\Gamma_{T}(A^{*}))}{\delta}\right)}{T}}\quad\text{ if }\quad T\gtrsim d\log\left(\frac{1+\operatorname{Tr}(\Gamma_{T}(A^{*}))}{\delta}\right).$ (1.2)
@@ 1.2 Learning multiple LDSs under smoothness constraints
which means that the quadratic variation of $A^{*}_{1},\dots,A^{*}_{m}$ w.r.t $G$ is bounded. For bounded $A^{*}_{l}$’s one can obviously find $S_{m}=O(\left|{\mathcal{E}}\right|)$ such that (1.5) is true, however the interesting smoothness regime is when $S_{m}=o(\left|{\mathcal{E}}\right|)$. Indeed, in this regime, the node-dynamics of the graph $G$ share a non-negligible amount of information between themselves, and thus one might hope to leverage this fact in order to jointly estimate $A^{*}_{l}$’s meaningfully. In this regard, the following “sanity checks” are useful.
@@
We propose two estimators for this problem, and derive non-asymptotic error bounds on the mean-squared error (MSE) $\frac{1}{m}\sum_{l=1}^{m}\big{\|}{\widehat{A}_{l}-A^{*}_{l}}\big{\|}_{F}^{2}$, where $\widehat{A}_{l}$ is the estimate of $A^{*}_{l}$. For both the estimators, we establish conditions under which the MSE goes to 00 as $m$ increases.
@@ 2.1 Notation
For any vector $x\in\mathbb{R}^{n}$, $\big{\|}{x}\big{\|}_{p}$ denotes the usual $\ell_{p}$ norm of $x$. For $X\in\mathbb{R}^{n\times m}$, we denote the spectral and Frobenius norms of $X$ by $\big{\|}{X}\big{\|}_{2}$ and $\big{\|}{X}\big{\|}_{F}$ respectively, while $\Big{\langle}X,Y\Big{\rangle}=\operatorname{Tr}(X^{\top}Y)$ denotes the inner product between $X$ and $Y$ (with $\operatorname{Tr}(\cdot)$ denoting the trace operator). For $n\times n$ matrices X with eigenvalues $\lambda_{i}\in\mathbb{C}$, we denote the spectral radius of $X$ by $\rho(X):=\max_{i=1,\dots,n}\left|{\lambda_{i}}\right|$. Also, $\kappa(X)(\geq 1)$ denotes the condition number of $X$ (ratio of largest and smallest singular values). For a positive definite matrix $B$, and a vector $x$, we will denote
@@
Finally, we denote the indicator function for an event $\mathcal{B}$ by $\mathbf{1}_{\mathcal{B}}$.
@@ 2.2 Setup
$\sum_{\left\{{l,l^{\prime}}\right\}\in\mathcal{E}}\big{\|}{A^{*}_{l}-A^{*}_{l^{\prime}}}\big{\|}_{F}^{2}={\color[rgb]{0,0,0}{(a^{*})}^{\top}}(L\otimes I_{d^{2}})a^{*}\leq S_{m}.$ (2.1)
@@ 2.3 Laplacian smoothing
The first estimator is the penalized least-squares (LS) method
$(\widehat{A}_{1},\dots,\widehat{A}_{m})=\underset{A_{1},\dots,A_{m}\in\mathbb{R}^{d\times d}}{\operatorname{argmin}}\ \left({\color[rgb]{0,0,0}\sum_{l=1}^{m}\sum_{t=1}^{T}\big{\|}{x_{l,{t+1}}-A_{l}x_{l,t}}\big{\|}_{2}^{2}}+\lambda\sum_{\left\{{l,l^{\prime}}\right\}\in\mathcal{E}}\big{\|}{A_{l}-A_{l^{\prime}}}\big{\|}_{F}^{2}\right)$ (2.3)
@@ 2.4 Subspace-constrained LS
Denoting $P^{(\tau)}=\sum_{l=m-\tau+1}^{m}v_{l}v_{l}^{\top}$, which is the projection matrix for the low-frequency part of $L$, we obtain another projection matrix $\widetilde{P}^{(\tau)}=P^{(\tau)}\otimes I_{d^{2}}$. The condition (2.1) implies that $a^{*}\in\mathbb{R}^{md^{2}}$ – formed by column-wise stacking of $\operatorname{vec}(A^{*}_{1}),\dots,\operatorname{vec}(A^{*}_{m})$ – lies close to the low frequency part of $L\otimes I_{d^{2}}$. Hence, $a^{*}\approx\widetilde{P}^{(\tau)}a^{*}$ for a suitable $\tau$ depending on $S_{m}$. For example, if $S_{m}=0$, then $\tau=1$ and clearly $a^{*}=\widetilde{P}^{(\tau)}a^{*}$.
@@ 3.1 Laplacian smoothing
The proof is motivated by ideas in the proof of (Sarkar and Rakhlin, 2019, Theorem 1) for learning marginally stable linear dynamical systems. The following points are useful to note.
  • The first error term in (1) corresponds to the bias (which increases as $\lambda$ increases), while the remaining two terms therein constitute the variance (which decrease as $\lambda$ increases). The bias-variance trade-off will be achieved by choosing $\lambda$ to minimize the sum of these terms.
  • The condition on $T$ in (3.1) is an artefact of the analysis, and arises from Lemma 1 in order to ensure $Q^{\top}Q+\lambda(L\otimes I_{d^{2}})\succ 0$. In general, obtaining lower bounds on the eigenvalues of $Q^{\top}Q+\lambda(L\otimes I_{d^{2}})$ is challenging as the eigenspaces of $L\otimes I_{d^{2}}$ and $Q^{\top}Q$ are not aligned. Our strategy is the admittedly naive approach that instead seeks to ensure $Q^{\top}Q\succ 0$ using results from (Sarkar and Rakhlin, 2019, Section 9).
@@
    The $T^{\alpha}$ term appearing in the numerator of the first term in (2) is essentially due to the $\gamma_{1}(T,m,\delta)$ term appearing outside the logarithm in the third term of (1). This is an artefact of the analysis, and it is unclear how to improve this for the case of marginally stable matrices. But when $A^{*}_{l}$’s satisfy a stricter stability condition where $\big{\|}{A^{*}_{l}}\big{\|}_{2}<1$ for each $l$, then this issue disappears. See Remark 3 below for details.
  2. 2.
@@
    Hence if $S_{m}=O(m^{1-\epsilon})$ for some constant $\epsilon\in(0,1)$, then we require $m=\Omega(\operatorname{poly}(d,T^{\alpha}))$ to drive the MSE below a given threshold. Recall that $\alpha\in[2d]$, and depends on the size of the largest Jordan blocks of $A^{*}_{l}$’s. If for each $l$, the largest Jordan block of $A^{*}_{l}$ is of constant size, then $\alpha\asymp 1$. But in case $\alpha\asymp d$, then $m$ will grow exponentially with $d$ in our bounds.
@@
Suppose each $A^{*}_{l}$ is marginally stable, i.e., satisfies (3.4), and recall $\alpha,\kappa_{\max}$ in (3.6). Under Assumptions 1 and 2, suppose $\eta_{l,t}$ has $R$-subgaussian entries for each $l,t$.
@@
    $\tilde{S}_{m}:=S_{m}+{\color[rgb]{0,0,0}m^{2}\Big{\|}A_{1}^{*}-\frac{1}{m-1}\sum_{l=2}^{m}A_{l}^{*}\Big{\|}_{F}^{2}}.$
@@
Let us interpret the above corollary.
  • The requirement $T=\Omega(d\log(\frac{dmT^{\alpha}}{\delta}))$ (for both complete and star graphs) is the same as that for a path graph. Also, the error bounds for both the examples contain the term $T^{\alpha}$ in the numerator (as in the case of the path graph) which again is an artefact of the analysis. As mentioned earlier, this issue disappears if $\big{\|}{A^{*}_{l}}\big{\|}_{2}<1$ for each $l$. See Remark 3 below.
  • When $G$ is the complete graph, the interesting regime for the smoothness is $S_{m}=o(m^{2})$. If $S_{m}=O(m^{2-\epsilon})$ for some $\epsilon$ in $(0,2)$, then we require $m=\Omega(\operatorname{poly}(d,T^{\alpha}))$ to drive the MSE below any given threshold. As in the case of a path graph, this bound will be exponential in $d$ if $\alpha\asymp d$.
  • When $G$ is the star graph, the term $\tilde{S}_{m}$ captures the smoothness of the problem. Not surprisingly, $\tilde{S}_{m}$ is small when a small fraction of the degree $1$ vertices have matrices which are far from $A^{*}_{1}$ (the matrix at the “central” node). If $\tilde{S}_{m}=O(m^{1-\epsilon})$ for $\epsilon$ in $(0,1)$, then we again require $m=\Omega(\operatorname{poly}(d,T^{\alpha}))$ (as in the other examples) to make the MSE small.
Remark 3 (Dependence on $T$ under stricter stability conditions).
The error bounds in Corollaries 1 and 2 (for the marginally stable case) involve the term $T^{\alpha}$ appearing in the numerator of the first term. As discussed earlier, this is an artefact of the analysis. However, in the relatively easier setting444which implies each $A^{*}_{l}$ is strictly stable as $\rho(A^{*}_{l})\leq\big{\|}{A^{*}_{l}}\big{\|}_{2}$. where each $A^{*}_{l}$ satisfies $\big{\|}{A^{*}_{l}}\big{\|}_{2}<1$, this issue disappears. Indeed, by noting that
${\color[rgb]{0,0,0}\operatorname{Tr}(\Gamma_{T}(A^{*}_{l}))=\sum_{k=0}^{T}\operatorname{Tr}(({A^{*}_{l}})^{k}(({A^{*}_{l}})^{k})^{\top})\leq\sum_{k=0}^{T}d\big{\|}{({A^{*}_{l}})^{k}(({A^{*}_{l}})^{k})^{\top}}\big{\|}_{2}\leq d\sum_{k=0}^{T}\big{\|}{A^{*}_{l}}\big{\|}_{2}^{2k}\leq\frac{d}{1-\big{\|}{A^{*}_{l}}\big{\|}_{2}^{2}}}$
and comparing it with (3.5), we can simply set
$\alpha=0\quad\text{and}\quad\kappa^{2}_{\max}=(1-\max_{l}\big{\|}{A^{*}_{l}}\big{\|}_{2}^{2})^{-1}$
in Corollaries 1 and 2 to obtain the error bounds in this setup. In particular, notice that the error bounds now go to zero as $T$ increases, which is of course what we would ideally like to have. Moreover, we now require $m=\Omega(\operatorname{poly}(d))$ to drive the MSE below a threshold.
@@
As a sanity check, let us verify that we obtain better bounds than what we would have by naively estimating each $A^{*}_{l}$ individually. Indeed, assuming $R$ to be a constant, if (3.4) holds, then applying the OLS for each node $l$, we obtain using (1.2), (3.5), and a simple union bound that with probability at least $1-\delta$,
$\frac{1}{m}\sum_{l=1}^{m}\big{\|}{\widehat{A}-A^{*}_{l}}\big{\|}_{F}^{2}\lesssim\frac{d^{2}}{mT}\sum_{l=1}^{m}\log\Big{(}\frac{m(1+\operatorname{Tr}(\Gamma_{T}(A^{*}_{l})))}{\delta}\Big{)}\lesssim{\color[rgb]{0,0,0}\frac{d^{2}}{T}\log\Big{(}\frac{mdT^{\alpha}\kappa_{\max}^{2}}{\delta}\Big{)}},$ (3.12)
provided $T\gtrsim d\max_{l\in[m]}\log(\frac{m[1+\operatorname{Tr}(\Gamma_{T}(A^{*}_{l}))]}{\delta})$. This condition on $T$ is essentially the same as in (3.1), and is ensured for $T=\Omega(d\log(\frac{dmT^{\alpha}}{\delta}))$. This latter condition on $T$ is ensured if $m\gtrsim dT^{\alpha}$ and $T=\Omega(d\log(m/\delta))$. However in the regime $T=\Theta(\log m)$, the bound in (3.12) is $\Omega(1)$ w.r.t $m$.